question is: solve the quadratic equation by completing the square:
x^2+6x=0
x^2+6x(6/2)^2 = (6/2)^2
(sqrt x + sqrt 3^2)^2 = sqrt 9
x+3= +- 3
x+3=3 x+3=-3
x=0 x=-6
This question is not from a textbook, I would like to know if I did the work correctly, thank you so much for your help.

Found 2 solutions by RAY100, solver91311:
Answer by RAY100(1637)

x^2 +6x =0
.
complete square by using (6/2)^2 as 3rd term,,,however must add to both sides to maintain balance.
.
x^2 +6x +(6/2)^2 = (6/2)^2
.
x^2 +6x +9 = 9
.
(x+3)^2 = 9
.
take square root of both sides
.
+/-(x+3) = +/-(3)
.
+ x+3 =+3,,,x=0,,,,check,,(0)^2 +6(0)=0,,,ok
.
-x-3 = -3,,,x=0,,,check,,ok
.
x+3=-3,,,x=-6,,check (-6)^2 +6(-6)=0,,,36-36=0,,,ok
.
-x-3=+3,,,,x=-6,,,,check,,,ok
.
Therefore,,,x=0,-6
.

Answer by solver91311(24713)


Your math notation is a little whacked, but as near as I can tell, you did the problem correctly. Just for future reference, it should have looked like:













or



John